### v2

# Toward Robust Semi-supervised Regression via Dual-stream Knowledge Distillation

Abstract
Semi-supervised regression (SSR), which aims to predict continuous scores of samples while reducing reliance on a large amount of labeled data, has recently received considerable attention across various applications, including computer vision, natural language processing, and audio and medical analysis. Existing SSR methods typically train models on scarce labeled data by introducing constraint-based regularization or ordinal ranking to reduce overfitting. However, these approaches fail to fully exploit the abundance of unlabeled samples. While consistency-driven pseudo-labeling methods attempt to incorporate unlabeled data, they are highly sensitive to pseudo-label quality and noisy predictions. To address these challenges, we introduce a Dual-stream Knowledge Distillation framework (DKD), which is specially designed for the SSR task to distill knowledge from both continuous-valued knowledge and distribution information, better preserving regression magnitude information and improving sample efficiency. Specifically, in DKD, the teacher is optimized solely with ground-truth labels for label distribution estimation, while the student learns from a mixture of real labels and teacher-generated pseudo targets on unlabeled data. The distillation design ensures the effective supervision transfer, allowing the student to leverage pseudo labels more robustly. Then, we introduce an advanced Decoupled Distribution Alignment (DDA) to align the target class and non-target class between teacher and student on the distribution, enhancing the student’s capacity to mitigate noise in pseudo-label supervision and learn a more well-calibrated regression predictor. Extensive experiments were conducted on diverse datasets, including audio, text, image, and medical data. We show that DKD exhibits strong generalization capabilities and outperforms the best competing methods by 4.97% and 11.81% in terms of overall MAE and $R^{2}$, respectively.
1 Introduction
Semi-supervised regression (SSR) aims to rank or predict continuous scores of samples while reducing reliance on a large amount of labeled data. It has been widely applied in various scenarios, including age estimation, clinical score prediction, and audio-based assessments, and other real-world applications where a large amount of labeled data across diverse are difficult to obtain Ren et al. (2022); Gui et al. (2024); Bao et al. (2024). With the rapid advancement of deep learning in recent years, regression tasks using deep learning models across various domains have attracted significant attention Yang et al. (2024); Van Engelen and Hoos (2020), especially in the SSR task. Some approaches directly employ deep learning models such as CNNs for images and RNNs for time series, optimizing them with loss functions like Mean Squared Error (MSE) or Mean Absolute Error (MAE) on the available labeled data Mohammadi Foumani et al. (2024); Yang et al. (2022); Zhang et al. (2017a). However, such a straightforward application struggles to learn effective patterns in scenarios with scarce labeled data, owing to the heavy reliance of deep learning models on large-scale annotations Pan et al. (2024); Yin et al. (2022). Therefore, direct regression (DR) based models typically yield suboptimal performance.
To exploit the abundance of unlabeled data, several consistency-driven pseudo-labeling methods that assign the pseudo-labels on the unlabeled samples have been proposed Zhang et al. (2025); Zheng et al. (2025). Pseudo-labels are typically generated by training on the available labeled samples, and consistency regularization is often used to assist the optimization process and pseudo-labeling Zhao et al. (2025); Zhao and Wen (2025). While pseudo-labeling helps mitigate the challenge of limited labeled data, as some predicted incorrect labels are directly treated as true labels during training, and the model is likely to overfit to the wrong annotations. Therefore, the effectiveness of these approaches heavily relies on the quality of the pseudo-labels Jo et al. (2024); Yin et al. (2022), and assessing the reliability of the generated pseudo-labels remains a challenging task.
Figure 1: MAE and test SRCC curves for direct regression (DR), Rankup, UCVME, and DKD on the BVCC dataset. Rankup yields the least competitive results, as its pairwise ranking method may not fully capture the continuous relationships. Although UCVME improves performance via consistency-constrained pseudo-labels, it still falls short of DKD.
To address these challenges, we introduce an advanced decoupled distillation framework (DKD), specifically designed for SSR, to effectively exploit unlabeled samples. DKD distills the knowledge from both continuous-valued knowledge and distribution information in an end-to-end manner, better preserving regression magnitude information and improving sample efficiency. Specifically, DKD is implemented by a teacher model optimized solely with ground-truth labels for label distribution learning, producing informative predictions, while the student learns from a mixture of real labels and teacher-generated pseudo scores on unlabeled data, ensuring effective supervision transfer. The distillation allows the student to leverage pseudo labels more robustly, mitigating the interference from noisy pseudo-supervision in SSR. We further introduce an advanced decoupling distribution alignment (DDA) approach to distill the distribution information from the labeled trained teacher model, where we decouple the distribution into two parts for alignment, rather than aligning the distribution directly. DDA aligns the target and non-target distributions between the teacher and student in a decoupled manner, with an adaptive learnable weighting scheme, thereby enabling the student to better exploit unlabeled samples while reducing overfitting to noisy pseudo-labels and learning a more well-calibrated regression model. As shown in the Fig 1, the DKD can learn more robust and more generalizable knowledge representation provided by the mixture of labeled and pseudo-labeled samples provided by the teacher model. It outperforms the conventional method, direct regression (DR), and strong competing methods, including RankUp Huang et al. (2024), UCVME Dai et al. (2023a), achieving the best performance on both MAE and SRCC. In summary, this work makes the following main contributions.
  • We propose DKD, the first end-to-end decoupled distillation framework for semi-supervised regression, which effectively directly distills both continuous-valued knowledge and distribution information in an end-to-end manner, better preserving regression magnitude information and improving sample efficiency under limited labels.
  • We further introduce a novel alignment method, block-aware decouple distribution alignment (DDA), which enables the student model to identify important classes, enhancing the student’s capacity to mitigate noise in pseudo-label supervision and learn a more well-calibrated regression predictor.
  • Extensive experiments were conducted on the datasets from various domains, including audio, text, image, and medical data. The experimental results demonstrate that DKD outperforms the state-of-the-art semi-supervised regression methods.
2 Related Works
2.1 Semi-supervised Regression
The SSR methods can be roughly categorized into pseudo-label-based methods, consistency regularization, and contrastive-learning-based methods. Jo et al. Jo et al. (2024) propose an SSR framework that uses uncertainty estimation to guide pseudo-labeling, alongside a pseudo-label calibration approach that enhances pseudo-label quality by propagating information from labeled to unlabeled samples. Semi-supervised deep kernel learning Jean et al. (2018) minimizes predictive variance on unlabeled data through consistency regularization, leveraging the robust uncertainty quantification provided by Gaussian processes. Some contrastive-learning-based methods exploit the ranking information among samples as auxiliary for regression Qiao et al. (2022); Huang et al. (2024). For example, RankUp Huang et al. (2024) constructs pairwise ranking labels from continuous targets with an auxiliary ranking classifier. CLSS Dai et al. (2023a) and GCLSS Wang et al. (2025) facilitate contrastive regression by exploiting the ordinal relationships between unlabeled samples with a spectral seriation algorithm.
Although these methods have achieved remarkable success in SSR, most of them do not fully leverage the potential of abundant unlabeled data. Pseudo-labeling approaches incorporate unlabeled samples into training, but they are often sensitive to noisy pseudo-targets and may even degrade performance. In DKD, we propose an advanced decoupled representation distillation strategy that transfers knowledge from both continuous-value supervision and distributional information, enabling effective utilization of both pseudo-labeled and ground-truth labeled samples.
2.2 Distribution Alignment in Distillation
Distribution alignment is a strategy used during student model training to match not only the outputs of the teacher model but also its underlying distributions. Existing approaches can be roughly categorized into logit-level alignment and feature-level alignment. Logit-level alignment aims to align the output distribution (e.g., softmax probabilities) of the teacher and student Hinton et al. (2015); Zhao et al. (2022); Zeng et al. (2025). For example, the multi-level logit distillation method Jin et al. (2023) employs multi-level prediction alignment to facilitate instance prediction learning in the student model. Curriculum temperature Li et al. (2023b) introduces temperature scheduling for dynamically adjusting softening intensity across training phases, which smooths gradient flow. The feature-level distribution alignment matches the representation, which is often done by maximum mean discrepancy, contrastive, and correlation-based loss Gou et al. (2021); Chen et al. (2021). Notable advancements in feature-level distillation encompass DiffKD Huang et al. (2023), training diffusion denoisers with teacher features to enforce attention map congruence in critical regions; WCoRD Chen et al. (2021), constructing global-local hybrid contrastive objectives with Wasserstein dual loss to align deep geometric structures. These alignment approaches have demonstrated effectiveness in knowledge transfer. Although there are some attempts by applying the decoupled distillation on the classification task Zhao et al. (2022); Li et al. (2023a); Feng et al. (2024). There is no work on semi-supervised regression that more emphasizes how pseudo-labels are leveraged effectively. Another difference of our method is that we incorporate the adaptive learnable weighting scheme specifically for the non-target alignment to further prevent the student from overfitting noisy pseudo-supervision. Meanwhile, the continuous-valued knowledge is incorporated into the distribution alignment on the basis of label distribution learning and alignment, to keep the regression magnitude information.
Figure 2: The Overview of DKD. (1) The input of DKD includes both labeled and unlabeled data. The teacher model is exclusively trained on the labeled data and is then used to generate pseudo labels for the unlabeled samples. The student model is trained on the entire dataset with a mixture of real labels and teacher-generated pseudo targets on unlabeled data. (2) Both the teacher and student are trained under a label-distribution formulation of regression, while continuous-valued distillation enforces consistency by minimizing the gap between their expected scores. (3) In the distribution distillation, DDA first identifies the target class to construct a binary probability distribution, then aligns the predicted distributions over classes between the teacher and student models separately for the target and non-target parts.
3 Methodology
Notations. We assume the data set $\mathcal{X}$ consists of $n$ labeled samples $\mathcal{X}_{l}=\{{({x}_{i},y_{i})}_{i=1}^{n}\}$ and $m$ unlabeled samples $\mathcal{X}_{u}=\{{({x}_{i})}_{i=m}^{n+m}\}$, with instance $x\in\mathbb{R}$ for image, text, audio or tabular medical data. $\mathcal{Y}=\{y_{i}\}$ denotes the set of ground truth labels, where each label $y_{i}$ is a continuous value in $\mathbb{R}$.
Problem Statement. Deep semi-supervised regression can be reformulated as given the training set $\mathcal{X}_{T}$ and $\mathcal{X}_{T}$ = $\mathcal{X}_{l}$, we aim to train a mapping function $f:\mathcal{X}_{u}\mapsto\mathcal{Y}$ which maps each input ${x}_{i}\in\mathcal{X}_{u}$ to a continuous scalar $y_{i}\in\mathbb{R}$. Semi-supervised learning explicitly incorporates unlabeled samples to complement limited labeled data during training. Therefore, in this case, the $\mathcal{X}_{T}$ consists of $\mathcal{X}_{l}$ and $\mathcal{X}_{u}$.
3.1 The Overview of DKD
As shown in Fig. 2, the proposed DKD predicts scores by establishing a decoupled representation distillation framework, where the ground truth is transformed into a label distribution, effectively converting the regression task into a discrete distribution estimation problem for both teacher and student models. We then apply DKD over the label distribution and DDA for the distribution alignment to enable the student model to learn fine-grained and more generalized knowledge.
This distillation process is particularly effective for the SSR task due to two key factors: (1) we reformulate the regression task as a distillation on both continuous-valued knowledge and distribution information in an end-to-end manner, which effectively mitigates the risk of the noise when integrating the pseudo labels, especially in the scenario that labeled samples are limited; and (2) we introduce a novel decoupled distribution alignment strategy for representation distillation over the classes, enabling the model to capture fine-grained features more effectively than conventional alignment methods. Finally, the model is jointly trained by combining the loss functions of the student and teacher, augmented with the proposed decoupled distribution alignment.
3.2 Distillation for Semi-supervised Regression
Label Distribution Learning Conventional SSR methods typically predict the score by directly mapping the input sample ${x}_{i}$ to the target value $y$. However, this approach is highly sensitive to noise and irrelevant features, often resulting in poor generalization and a high risk of overfitting, especially when few labels are available. Therefore, following Gao et al. (2017); Xu et al. (2019), we define several classes $\mathcal{B}=\{{\mathbf{b}_{i}}\}_{i=1}^{L}$ based on the range of ground truth $y$, converting the regression task into a discrete distribution estimation problem. $\mathcal{B}$ can be intuitively viewed as a set of evenly spaced score bins (anchors) that discretize the continuous regression label range into $L$ classes, each covering an interval of width $c$, $\mathbf{b}_{i}=(i-1)\times c$, where each score is assigned to one corresponding class. Then, $y$ can be obtained by calculating the expectation as follows:
$y=\sum\limits_{i=1}^{L}{{\tilde{p}_{i}}}{\mathbf{b}_{i}}$ (1)
where $\tilde{p}_{i}$ is the output probability of the model with the corresponding class $\mathbf{b}_{i}$}. Therefore, the mapping function $f$ in SSR is modified to learn a distribution $\mathcal{P}=\{p_{i}\}_{i=1}^{L}$ over predefined classes. By transforming the numerical regression into an interval prediction problem, it offers a smoother learning objective and richer label distribution than direct numerical regression Xu et al. (2019), leading to more effective use of limited labels and better noise robustness.
To obtain a probabilistic distribution, we apply the softmax function, and the predicted distribution is formulated as:
$p_{i}=\frac{\exp\left(\tilde{p}_{i}\right)}{\sum_{i=1}^{L}\exp\left(\tilde{p}_{j}\right)}$ (2)
where $p_{i}$ is the predicted probability for the class $\mathbf{b}_{i}$ after softmax. These outputs form a distribution, from which we compute the expectation to obtain the final score as follows:
$\hat{y}_{i}=\sum\limits_{i=1}^{L}{{p_{i}}}{\mathbf{b}_{i}}$ (3)
Finally, we employ the MAE loss to minimize the distance between the prediction $\hat{y}_{i}$ and the ground truth $y_{i}$:
$\mathcal{L}=\frac{1}{|{{\cal X}}|}\sum_{x_{i}\in{\cal{X}}}|y_{i}-\hat{y}_{i}|$ (4)
where $y_{i}$ and the $\hat{y}_{i}$ are the ground truth label and the predicted score of sample $x_{i}$ , respectively. $|{\cal X}|$ is the number of samples in $\cal X$}. In DKD, both the teacher and student models utilize the MAE loss function to incorporate supervised information, as detailed below.
Continuous-valued Knowledge Distillation Since using label distribution learning alone is insufficient for effective representation learning, we reformulate the knowledge distillation process to enable the student model to learn more generalized features for improved performance in SSR. To build the distillation framework for SSR, we first apply weak augmentation like random crop, flip, to the data before feeding it into the teacher model, while strong augmentation, like RandAugment Cubuk et al. (2020) is applied to the input of the student model. The detailed description of strong/weak augmentation on different domain data can be found in App. C.
The predicted scores from the teacher model are used to guide the student by assigning corresponding pseudo-labels to the unlabeled nodes. As a result, the student model is trained on the entire dataset, including both labeled samples and unlabeled samples with pseudo-labels. The loss functions for the teacher and student models are thus defined separately as follows.
$\mathcal{L}_{\mathcal{T}}=\frac{1}{|{{\cal X}_{l}}|}\sum_{x_{i}\in{\cal{X}}_{l}}|y_{i}-\hat{y}_{i}^{\mathcal{T}}|$ (5)
$\mathcal{L}_{\mathcal{S}}=\frac{1}{|{{\cal X}}|}\sum_{x_{i}\in{\cal{X}}}|y_{i}-\hat{y}_{i}^{\mathcal{S}}|$ (6)
where $\hat{y}_{i}^{\mathcal{T}}$ and $\hat{y}_{i}^{\mathcal{S}}$ are the predicted scores of the teacher model and the student model, respectively. To enable the student model to learn more accurate information for SSR, we define the loss $\mathcal{L}_{\text{DKD}}$ as follows, based on the joint training of the teacher and student models.
$\mathcal{L}_{DKD}=\mathcal{L}_{\mathcal{T}}+\mathcal{L}_{\mathcal{S}}+\mathcal{L}_{Aligment}$ (7)
The joint training mechanism enables the teacher model to generate high-quality pseudo labels from a small amount of labeled data, continuously improving the utilization of unlabeled data. The alignment loss $\mathcal{L}_{Alignment}$ minimizes the distance between the representations of the student and teacher models, typically implemented using Kullback–Leibler (KL) divergence or MSE. In DKD, we align the label distributions and further introduce a decoupled distribution alignment strategy, which will be described in the following sections.
3.3 Decoupling Distribution Alignment (DDA)
Conventional alignment methods often simply use KL divergence or MSE to minimize the distance between the teacher and student models. Although distillation helps the student exploit pseudo-labels in a more stable manner, simple alignment may be misled by noisy targets in pseudo-labeling data, yielding suboptimal performance Mei et al. (2023). To mitigate this issue, we propose a DDA strategy that aligns the distributions over the classes at a more granular level, improving the student’s ability to fully exploit unlabeled samples and learn a better-calibrated regression model Besides, aligning the distribution of discretized prediction classes is essential for effectively transferring knowledge from the teacher to the student. Furthermore, we empirically demonstrate that employing more advanced alignment strategies enables the student model to acquire richer and more effective knowledge, thereby enhancing the generalization ability of DKD.
Inspired by the success of decoupling target-level supervision from non-target relational insights in the classification task Zhao et al. (2022); Li et al. (2024); Zheng and Cheng (2025), we adopt a similar strategy to better exploit latent information from unlabeled samples in our SSR framework. Specifically, we first identify the target classes to formulate a binary classification task, treating the remaining classes as non-targets. The alignment is then performed separately on the target and non-target classes.
Target class identification. Since the entire dataset is fed into the teacher model, it is worth noting that the unlabeled samples do not provide supervision signals—only the labeled samples contribute to the MAE loss in the teacher model. But in the advanced alignment, we choose to involve the unlabeled samples to provide more consistent information. Specifically, to determine the target class, we utilize the ground-truth labels for the labeled samples and the pseudo-labels for the unlabeled ones. In the presence of ground-truth labels, the target class $\mathbf{b}_{t}$ is assigned to the corresponding label $y_{i}$; otherwise, $\mathbf{b}_{t}$ can be derived from the teacher’s prediction $\hat{y}_{i}^{\mathcal{T}}$. Furthermore, we denote the probability associated with the target class $\mathbf{b}_{t}$ as $p_{t}$. The probability of the non-target set $\mathcal{B}\setminus\{\mathbf{b}_{t}\}$ is thus defined as $1-p_{t}$.
Target and non-target class distribution alignment. Based on the identification of the target class, we can split the class into the target distribution to formulate the $\mathbf{p}={\{p_{t},1-p_{t}\}}$, with $\mathbf{p}^{\mathcal{T}}$ and $\mathbf{p^{\mathcal{S}}}$ to represent the teacher and student. The results of classes refer to the non-target distribution, i.e., the remaining classes after excluding the target class, denoted as $\mathbf{q}=\{p_{k}\}_{k\neq t}^{L}$, with $\mathbf{q}^{\mathcal{T}}$ and $\mathbf{q^{\mathcal{S}}}$ to represent the teacher and student. The optimization of DDA can be reformulated as follows:
$\mathcal{L}_{DDA}=\frac{1}{|{{\cal X}}|}\sum_{x_{i}\in{\cal{X}}}\mathrm{KL}\left(\mathbf{p}_{i}^{\mathcal{T}},\mathbf{p}_{i}^{\mathcal{S}}\right)+\beta\cdot\sum_{x_{i}\in{\cal{X}}}\omega_{i}\cdot\mathrm{KL}\left(\mathbf{q}_{i}^{\mathcal{T}},\mathbf{q}_{i}^{\mathcal{S}}\right)$ (8)
where $\mathrm{KL}\left(\mathbf{p}_{i}^{\mathcal{T}}\|\mathbf{p}_{i}^{\mathcal{S}}\right)$ aims to measure the similarity between the teacher distribution and student distribution on the target class, $\mathrm{KL}\left(\mathbf{q}_{i}^{\mathcal{T}}\|\mathbf{q}_{i}^{\mathcal{S}}\right)$ is the similarity on the non-target class. $\beta$ is the hyperparameter that controls the weights of no-target class alignment. For each sample, we introduce one learnable weight $\omega_{i}$ which is defined as $max(1-\sigma_{i}^{\mathcal{T}}$, 0), where $\sigma^{\cal{T}}_{i}$ is the standard deviation of teacher predictions. The key intuition is that incorporating this variance-aware weighting into the non-target class alignment allows the model to down-weight uncertain samples and focus the alignment on more reliable teacher guidance.
Compared with directly minimizing the KL divergence between teacher and student predictions on a mixture of labeled and pseudo-labeled samples, DDA performs decoupled alignment on target and non-target distributions, and introduces an adaptive learnable weighting scheme specifically for the non-target alignment. This prevents the student from overfitting noisy pseudo-supervision. As a result, DDA provides a more robust distribution alignment, enabling the student to better exploit unlabeled data under mixed supervision of ground-truth labels and pseudo targets.
3.4 Training and Inference
Training. During training, the DKD adopts the proposed distillation framework by jointly training the teacher and student model with corresponding loss functions $\mathcal{L}_{T}$ and $\mathcal{L}_{S}$. The $\mathcal{L}_{Alignment}$ is implemented by the $\mathcal{L}_{DDA}$, defined as follows:
$\mathcal{L}_{DKD}=\mathcal{L}_{T}+\mathcal{L}_{S}+\mathcal{L}_{DDA}$ (9)
Inference. During inference, we solely use the student model to predict the score of a sample ${x}_{i}$ in the test set $\mathcal{X}_{Test}$, as the student model is expected to learn rich and generalized knowledge from the teacher by leveraging the DDA. Give the unlabeled sample ${x}_{i}$ in the $\mathcal{X}_{Test}$, the predicted score for the unlabeled sample, denoted as $\hat{y}_{i}$:
$\hat{y}_{i}=\sum\limits_{i=1}^{L}{{p_{i}^{\mathcal{S}}}}{\mathbf{b}_{i}}$ (10)
where $p_{i}^{\mathcal{S}}$} is the prediction of the student model and $\mathbf{b}_{i}$ is the value of the class.
4 Experiments
Datasets. We conduct experiments on four real-world regression benchmark datasets spanning computer vision, speech, natural language processing, and clinical data analysis: (1) BVCC Cooper and Yamagishi (2021) is an audio dataset that is utilized for predicting the perceptual quality of audio samples. (2) Yelp Asghar (2016) is a textual benchmark of customer reviews used to predict sentiment intensity. (3) UTKFace Zhang et al. (2017b) is an image dataset used to estimate age from facial images. (4) MIMIC Johnson et al. (2023) is a medical tabular dataset that is used to predict the Sequential Organ Failure Assessment (SOFA) score. Table 1 summarizes their key statistics. More details about datasets are available in APP. A.
Model BVCC Yelp UTKFace MIMIC
# Train Set 4,974 250,000 18,964 3,623,503
# Test Set 1,066 25,000 4,741 55,859
Domain Audio Text Image Medical
Table 1: The statistical information of four datasets
Model BVCC Yelp UTKFace MIMIC
MAE $\downarrow$ $R^{2}$ $\uparrow$ SRCC $\uparrow$ MAE $\downarrow$ $R^{2}$ $\uparrow$ SRCC $\uparrow$ MAE $\downarrow$ $R^{2}$ $\uparrow$ SRCC $\uparrow$ MAE $\downarrow$ $R^{2}$ $\uparrow$ SRCC $\uparrow$
DR 0.533 0.490 0.741 0.723 0.566 0.769 9.420 0.540 0.712 2.726 0.239 0.510
$\pi$-Model (2017) 0.534 0.489 0.740 0.730 0.565 0.769 9.450 0.534 0.706 2.725 0.240 0.506
Mean Teacher (2017) 0.532 0.492 0.742 0.730 0.565 0.769 8.850 0.586 0.745 2.724 0.240 0.507
MixMatch (2019) 0.597 0.353 0.626 0.886 0.381 0.660 7.950 0.692 0.832 2.707 0.252 0.514
CLSS (2023) 0.499 0.534 0.748 0.721 0.543 0.748 9.100 0.586 0.737 2.714 0.251 0.510
UCVME (2023) 0.498 0.553 0.774 0.775 0.540 0.763 8.630 0.626 0.767 2.656 0.285 0.539
RankUp (2024) 0.470 0.588 0.776 0.661 0.645 0.829 7.060 0.751 0.835 2.720 0.244 0.509
GCLSS (2025) 0.455 0.613 0.786 0.592 0.669 0.821 7.018 0.743 0.829 2.934 0.135 0.374
DKD (Ours) 0.422 0.669 0.818 0.549 0.694 0.839 6.852 0.761 0.838 2.629 0.291 0.541
Table 2: The comparison results on four datasets. The best performance is boldfaced, with the second-best underlined.
Competing Methods. The competing methods can be categorized into three groups: direct regression, pseudo-label generation, and contrastive-learning-based approaches. (1) We directly predict the score by optimizing it using the MAE loss function, referred to as Direct Regression (DR). (2) The pseudo-label generation-based methods typically leverage the consistency regularization methods, including five technical approaches: perturbation-based self-ensembling frameworks, including $\pi$-model Laine and Aila (2017); momentum-averaged teacher mechanisms with exponentially moving average, such as mean teacher Tarvainen and Valpola (2017); co-training architectures like the dual-branch framework UCVME Dai et al. (2023a); entropy-regularized constraint approaches typified by MixMatch Berthelot et al. (2019); (3) The contrastive-learning-based methods leverage the ordinal relationships between unlabeled samples, including: the instance-level contrastive approach like CLSS Dai et al. (2023a); GCLSS Wang et al. (2025) extends CLSS by leveraging a mixed similarity matrix from both labeled and unlabeled samples; RankUp Huang et al. (2024) improves SSR with an auxiliary ranking classifier. More details can be found in App. B.
Evaluation Metric. Following the previous work Huang et al. (2024); Dai et al. (2023a), we adopt three widely used metrics to quantitatively assess the performance of DKD and other competing methods, including MAE, the correlation coefficient ($R^{2}$), and the Spearman Rank Correlation Coefficient (SRCC). MAE quantifies the average absolute deviation between the model’s predictions and the true values. Lower MAE indicates better predictive accuracy. The SRCC quantifies the monotonic relationship between the predicted and ground-truth rankings, reflecting the proportion of variance in the target variable captured by the model.
Model BVCC Yelp UTKFace MIMIC
[…]
S-LDL 0.441 0.640 0.795 0.717 0.514 0.728 9.570 0.505 0.687 2.748 0.227 0.481
DKD-Logits 0.506 0.529 0.737 0.576 0.655 0.819 7.981 0.648 0.799 2.707 0.248 0.520
DKD-KL 0.500 0.547 0.747 0.610 0.617 0.8033 7.502 0.693 0.808 2.674 0.264 0.531
DKD (Ours) 0.422 0.669 0.818 0.549 0.694 0.839 6.852 0.761 0.838 2.629 0.291 0.541
Table 3: Ablation studies on the proposed model. The best performance is highlighted in bold, with the second-best underlined.
Implementation Details. Our DKD model is implemented in Python 3.10 with PyTorch 2.1.2. All experiments are conducted on a GeForce RTX 4060 GPU with 8 GB of memory. The experimental configurations for DKD strictly adhere to those described in Huang et al. (2024). In this paper, we set the default value of $\beta$ to 10. We also adjust the $\beta$ to explore its effectiveness, see Fig. 4. For all of the datasets. The number of classes $L$ is set to 200 by default for the BVCC, Yelp, and UTKFace datasets and 100 for the MIMIC dataset. To ensure fair comparison across methods, we adopt the same augmentation schemes in RankUp Huang et al. (2024). We apply weak augmentation to enrich the teacher’s knowledge, while strong augmentation is applied to the student to enhance generalization. More detailed parameter configurations and data augmentation mechanisms are provided in the App. C.
4.1 Main Comparison Results
Table 2 summarizes the performance of various methods across four datasets from four domains. All semi-supervised methods were trained on 250 labeled samples and evaluated on the test set. From the results, we observe that DKD significantly outperforms the tested SSR methods across all datasets. Specifically, DKD outperforms the second-best model, GCLSS, achieving significant overall improvements of 4.97% in MAE, 11.8% in $R^{2}$, and 8.04% in SRCC across all datasets. Notably, GCLSS suffers a significant performance degradation on the MIMIC dataset. The possible reason is that GCLSS heavily relies on the feature similarity matrix, which is sensitive to the noise inherent in the MIMIC dataset derived from real-world ICU environments. Conversely, DKD achieves SOTA performance by fully leveraging unlabeled samples to learn more robust and generalizable knowledge representations through DDA.
The consistency regularization-based methods like the $\pi$-model, Mean Teacher, and MixMatch, which apply a consistency loss based on various data augmentations, consistently exhibit inferior performance across most datasets, merely matching the efficacy of DR on most datasets. The main reason is that their hybrid augmentation struggles to generate high-quality numeric pseudo-labels for the regression task, particularly under extremely limited sample conditions. In contrast, DKD leverages decoupled distillation, thereby not only enhancing pseudo-label generation quality but also more effectively utilizing target and non-target distributions, which collectively improve its suitability for SSR tasks.
As for the co-trained method, UCVME, which maximizes variational mutual information, yields suboptimal results on the MIMIC dataset. However, it requires a more complex Bayesian neural network to perform uncertainty variational inference for real-valued pseudo-label refinement. In contrast, DKD distills knowledge from both continuous-valued knowledge and distribution information, consequently surpassing UCVME. Moreover, DKD also consistently outperforms CLSS by learning more generalized feature representations through superior distribution alignment.
4.2 Ablation Study
In this section, we perform an ablation study to evaluate the contribution of each component in DKD. We design three variants, including (1) Single-Label Distribution Learning (S-LDL), we remove the distillation framework, only keep the one branch, and train on the labeled samples. This is to evaluate the effectiveness of the proposed distillation framework. (2) DKD-Logits, which remove the DDA i.e., only label alignment. (3) DKD-KL, which replaces the DDA with KL divergence. Both DKD-KL and DKD-Logits are used to evaluate the effectiveness of DDA. The results are shown in Table 3. From the results, the DKD can consistently outperform all the variants and yields the best performance on all the datasets, demonstrating its effectiveness. Both DKD-KL and DKD-Logits underperform the full DKD model but outperform S-LDL on the Yelp, UTKFace, and MIMIC datasets. This demonstrates that distillation effectively captures generalized patterns across text, image, and tabular medical data modalities. On the BCVV dataset, S-LDL achieves the second-best performance, though still behind DKD. This suggests that while S-LDL, leveraging only label distribution learning, can effectively learn patterns in audio data, the absence of distillation and DDA limits its ability, leading to suboptimal results.
Figure 3: The MAE values w.r.t of labeled sample size $n$.
Figure 4: The effectiveness of hyperparameters $L$ and $\beta$.
4.3 Performance w.r.t. Labeled Sample Size
To explore the impact of labeled samples, we compare DKD with the competing methods, using varying numbers of labeled samples, with the results reported in Fig. 3. As the number of labeled instances $n$ grows, DKD generally improves across the datasets, which is consistent with other tested semi-supervised methods. Across various labeling ratios, our DKD consistently achieves the best performance with the lowest MAE, demonstrating the robustness and efficacy of our proposed method.
4.4 Sensitive Analysis
We evaluate the sensitivity of DKD w.r.t. the hyperparameter class size $L$ and $\beta$ in $\mathcal{L}_{DDA}$. The results are shown in Fig. 4. More results can be found in App.D.
Performance w.r.t. Hyperparameter $L$: As shown in the Fig. 4(a), DKD performs stably under different classes on text, image, and audio datasets, and the performance slightly improves as $L$ increases. The main reason lies in the capacity of larger classes to capture the intricate relationships within the label distribution more accurately. In contrast, on the MIMC dataset, DKD achieved its highest performance with 100 classes, primarily because SOFA scores are more concentrated within a specific interval (typically around 4-10 in certain cohorts) compared to other datasets, and smaller classes are capable of capturing sufficient label information.
Performance w.r.t. Hyperparameter $\beta$ in $\mathcal{L}_{DDA}$: As shown in the Fig. 4(b), DKD exhibited a consistent trend across all tested datasets, achieving optimal performance when beta was set to 10. The main reason is that a smaller $\beta$ might cause the model to emphasize the ’difficulty’ of training samples, while a larger $\beta$ forces it to learn more generalized non-target information from the student network. An effective balance is achieved when beta is set to $10$.
4.5 Efficiency Analysis
Time Complexity In this section, we analyze the time complexity of DKD. Given that both models share the same backbone output as input, the computational cost of the backbone is excluded from this analysis. (1) The label distribution learning module comprises a fully connected layer and an expectation computation. Its time complexity is $\mathcal{O}(LM+2L)$, where $M$ denotes the dimension of the backbone output feature and $L$ represents the number of classes. (2) The DKD consists of two components, including DDA and continuous-valued knowledge distillation. Their respective time complexities are $\mathcal{O}(L)$ and $\mathcal{O}(1)$ per sample. Consequently, the overall time complexity of DKD is $\mathcal{O}(N(2LM+5L))$, where $N$ is the number of samples.
Model MIMIC Yelp UTKFace BVCC
UCVME (2023) 19 108 30 112
RankUp (2024) 10 28 43 46
GCLSS (2025) 8 23 12 36
DKD (ours) 15 34 150 66
Table 4: Runtimes (in seconds) on the four datasets.
Running Time Analysis. The runtimes of DKD and two recently proposed semi-supervised methods, including both training and inference time on NVIDIA GeForce GTX 4060, are reported in Table 4. DKD demonstrates higher efficiency on tabular medical data, text, and audio datasets. Although it requires more time on image datasets, this overhead is acceptable given the performance gains. Overall, DKD offers a favorable trade-off between efficiency and effectiveness.
5 Conclusion
In this paper, we propose DKD, the very first dual-stream knowledge distillation framework for fully exploiting the abundance of unlabeled samples and noise mitigation in pseudo-label supervision in SSR. DKD is designed to distill both continuous-valued knowledge and distribution information, which better preserves regression magnitudes and enhances the sample efficiency. Then, the DDA is proposed to further enhance the student’s capacity to mitigate noise in pseudo-label supervision and learn a more well-calibrated regression predictor. Extensive experimental results across multiple datasets spanning four domains demonstrate that DKD outperforms state-of-the-art methods, indicating its superior generalization capability.
Appendix A Datasets
A detailed introduction of all the datasets used in our work is given as follows.
BVCC Cooper and Yamagishi [2021] serves as an audio quality assessment benchmark, where the objective is to predict the perceptual quality of audio samples. Annotated on a 1-to-5 scale, the labels represent averaged scores from multiple listeners. This dataset comprises 4,974 training samples, 1,066 validation samples, and 1,066 test samples. Following the protocol defined in Rankup Huang et al. [2024], we exclusively utilize the training and validation splits for performance evaluation.
Yelp Asghar [2016] focuses on textual opinion mining, targeting the prediction of customer ratings based on reviews posted on the Yelp platform. The five-class rating labels represent distinct satisfaction levels. We employ the preprocessed Yelp data provided in the USB codebase Huang et al. [2024], consisting of 250,000 training samples, 25,000 validation samples, and 10,000 test samples. Consistent with RankUp’s methodology, only the training split is used for model training, while evaluation is performed on the validation set.
UTKFace Zhang et al. [2017b] addresses image-based age estimation, aiming to predict the age of individuals depicted in facial images. The age labels range from 1 to 116 years. RankUp partitions this dataset into 18,964 training samples and 4,741 test samples, with experiments conducted on its aligned and cropped version.
MIMIC Johnson et al. [2023] presents a multivariate time-series score prediction task, where the objective is to forecast SOFA scores (indicating illness severity from 0 to 24) based on patients’ clinical features over time. This dataset contains 3,623,503 training samples and 55,859 test samples.
Models
Wide
ResNet-28-2
Whisper
Base
Bert
Small
Bert
Training Iterations 262,144 102,400 102,400 1024,000
Evaluation Iterations 1,024 1,024 1,024 1,024
Batch Size 32 8 8 8
Optimizer SGD AdamW AdamW AdamW
Momentum 0.9 - - -
Criterion MAE MAE MAE MAE
Weight Decay 1e-03 2e-05 5e-04 2e-5
Layer Decay 1.0 0.75 0.75 0.75
Learning Rate 1e-02 2e-06 1e-05 2e-06
EMA Weight 0.999 - - 0.999
Pretrained False True True False
Sampler Random Random Random Random
Image Resize 40x40 - - -
Max Length Seconds - 6.0 - -
Sample Rate - 16,000 - -
Max Length - - 512 -
Table 5: The default hyperparameters for the base models.
Models $\pi$-Model MeanTeacher MixMatch UCVME CLSS GCLSS RankUp DKD
Unlabeled Batch Ratio 1.0 1.0 1.0 1.0 0.25 0.25 7.0 7.0
Regression Unlabeled Loss Ratio 0.1 0.1 0.1 0.05 - - - -
Regression Unlabeled Loss Warmup 0.4 0.4 0.4 - - - - -
Mixup Alpha - - 0.5 - - - - -
Dropout Rate - - - 0.05 - - - -
Ensemble Number - - - 5 - - - -
CLSS Lambda - - - - 2.0 2.0 - -
Labeled Contrastive Loss - - - - 1.0 1.0 - -
Unlabeled Contrastive Loss - - - - 0.05 0.05 - -
Unlabeled Rank Loss Ratio - - - - 0.01 0.01 - -
ARC Unlabeled Loss Ratio - - - - - - 1.0 -
ARC Loss Ratio - - - - - - 0.2 -
Confidence Threshold - - - - - - 0.95 -
Temperature - - - - - - 0.5 -
The Number of Selected Features - - - - - 7 - -
$\beta$ - - - - - - - 10
Table 6: Specific hyperparameters for each semi-supervised regression method.
Appendix B Competing Methods
A more detailed introduction of the six competing models and the evaluation metrics adopted in our work is given as follows.
B.1 Description of Baselines
$\pi$-Model Laine and Aila [2017] achieves self-ensembling by aggregating the outputs of a single network at different training stages. Consensus predictions for unknown labels are formed through diverse regularization and augmentation strategies. Perturbations are injected into both branches, and consistency between the two perturbed outputs is enforced, ensuring a more balanced and symmetric comparison.
Mean Teacher Tarvainen and Valpola [2017] aims to train a model via consistency between original samples and their perturbed counterparts, while using an exponential-moving-average (EMA) version of the model—termed the teacher—to supply more reliable guidance to the current student during training.
MixMatch Berthelot et al. [2019] consolidates the dominant approaches in semi-supervised learning by generating low-entropy label guesses for augmented unlabeled examples and subsequently blending labeled and unlabeled data via MixUp.
UCVME Dai et al. [2023b] improves training by producing high-quality uncertainty estimates for pseudo-labels via heteroscedastic regression. By explicitly modeling label uncertainty, the approach assigns greater importance to more reliable pseudo-labels. Furthermore, a novel variational model-ensembling scheme is introduced to reduce predictive noise and generate stronger pseudo-labels.
CLSS Dai et al. [2023a] extends contrastive regression to the semi-supervised domain. It exploits a spectral ordering algorithm to extract ordinal rankings from the feature-similarity matrix of unlabeled data and then leverages these rankings as supervisory signals. The additional ranking loss enhances robustness, marking the first exploration of contrastive learning for regression tasks.
RankUp Huang et al. [2024] reframes the original regression task as a ranking problem and trains it jointly with the regression objective. By introducing an auxiliary ranking classifier whose outputs can be handled by any off-the-shelf semi-supervised classification method, RankUp seamlessly integrates existing classification techniques into regression-oriented semi-supervised learning.
GCLSS Wang et al. [2025] extended CLSS to allow both labeled and unlabeled data to be used in the semi-supervised setting. Particularly, GCLSS uses labeled samples to regularize and rectify the recovered ordinal rankings of unlabeled data. The recovered rankings serve as pseudo-labels to supervise both the feature representations and the final predictions.
Dataset Augmentation
UTKFace Weak: Random Crop, Random Horizontal Flip
Strong: RandAugment Cubuk et al. [2020]
BVCC Weak: Random Sub-sample
Strong: Random Sub-sample, Mask, Trim, Padding
Yelp Weak: None
Strong: Back-Translation Xie et al. [2020]
MIMIC Weak: Random Masking
Strong: Operation Group
Table 7: Data augmentation strategies for different datasets.
B.2 Evaluation Metric
We used Mean Absolute Error (MAE), $R^{2}$, and Spearman’s Rank Correlation Coefficient (SRCC) as evaluation metrics for the tested model. The calculation process for each metric is as follows:
MAE: measures the average magnitude of the errors in a set of predictions, defined as follows:
$\mathrm{MAE}=\frac{1}{N}\sum_{i=1}^{N}|\hat{y}_{i}-y_{i}|,$
where $N$ denotes the number of samples, $\hat{y}_{i}$ is the model’s prediction, and $y_{i}$ is the corresponding ground-truth.
$\bf{R^{2}}$ measures the proportion of variance in the ground-truth labels that is captured by the model. It is defined as
$R^{2}=1-\frac{\sum_{i=1}^{N}(\hat{y}_{i}-y_{i})^{2}}{\sum_{i=1}^{N}(\bar{y}-y_{i})^{2}},$
where $\bar{y}$ is the mean of the ground-truth labels. $R^{2}$ ranges from $(-\infty,1]$; higher values indicate superior explanatory power. Let r(y^)r($\hat${y}) and $r(y)$ denote the ranks of predictions and labels, respectively. Then
SRCC is a non-parametric measure of the statistical dependence between the rankings of two variables.
$\mathrm{SRCC}=\frac{\mathrm{cov}(r(\hat{y}),r(y))}{\sigma_{r(\hat{y})}\,\sigma_{r(y)}},$
with $\mathrm{cov}(\cdot,\cdot)$ and $\sigma$ denoting covariance and standard deviation, respectively. SRCC ranges from [1,1][-1,1], where the value closer to $1$ implies a stronger rank-order.
Appendix C Parameters and Data Augementation
C.1 Parameter Configurations of Base Models
In this work, we employ a distinct base model for each database: Whisper-Base for audio data on BVCC, Bert-Small for text data on Yelp, Wide ResNet-28-2 for image data on UTKFace, and BERT for time-series prediction of SOFA values on MIMIC medical data. The common parameter Configurations of all the base models as listed in Table 5. Specific hyperparameter configurations for each SSR method are provided in Table 6.
C.2 Weak and Strong Augmentation
For audio, image, and text data, we adopt the universal semi-supervised learning benchmark framework, USB Wang et al. [2022], to implement the weak and strong augmentations following the previous work Huang et al. [2024]. For the MIMIC dataset, given its time-series nature, weak augmentation involves applying random masking operations, while strong augmentation employs $m$ rounds of random masking combined with Gaussian noise injection for multivariate time-series data. Specifically, each random masking operation obscures 5% of the observed data. Each Gaussian noise application introduces additive noise sampled from a normal distribution with zero mean and a variance of 0.02. To ensure fair comparison across methods, the same augmentation strategy is applied to both DKD and other comparative methods.
Detailed descriptions of data augmentation methods across different domains are presented in Table 7.
(a) The MAE w.r.t of bukets sise $L$.
(b) The MAE w.r.t of $\beta$.
Figure 5: The effectiveness of the hyperparameters.
Appendix D Additional Experimental Results
In this section, we further evaluate the effectiveness of the hyperparameters, specifically the class size of $L$ and $\beta$, in $\mathcal{L}_{DDA}$ using additional evaluation metrics. The results are shown in Fig.5. For the effectiveness of the class size $L$, compared to the $R^{2}$, the MAE for DKD exhibits greater stability under variations of $L$, see Fig.5 (a). We found that the best performance is achieved with 200 classes for the BVCC, Yelp, and UTKFace datasets, while a lower MAE is obtained for the MIMIC dataset using 100 classes. Regarding the effectiveness of $\beta$ in $\mathcal{L}_{DDA}$, see Fig.5 (b), the MAE demonstrates a trend comparable to that of $R^{2}$ across all datasets, with optimal performance universally observed at $\beta=10$.
(a) A prediction example of DKD with ground truth $y=4$.
(b) A prediction example of DKD with $y=2$.
Figure 6: The visualization results of DKD.
Appendix E Visualization Results
Fig. 6 presents a visualization of the prediction cases using the DKD model, illustrating the predicted discrete classes and their associated probability values on the Yelp dataset. As shown in Fig. 6 (a), for a regression value of 2, the Yelp dataset score prediction is made using a distributed set of LL=200 classes that span the 0-to-5 rating range. Notably, in addition to the score of 2 (Fair/Dissatisfied) itself, the model primarily leverages the probabilities of the nearby related values of 1 (Poor/Very Dissatisfied) and 3 (Average/Neutral) for a combined prediction. A similar phenomenon can also be observed in the other example in Fig 6 (b). These results highlight DKD’s superior ability to exploit label information, which significantly enhances generalization and prevents overfitting.